OpenStudy (anonymous):

Solve the system of equations. 2x+3y=1 y=3x+15 (?.?)

OpenStudy (solomonzelman):

the way they are setting the system up they are offering substitution. you have a system \(\large\color{blue}{ \displaystyle 2x+3y=1}\) \(\large\color{blue}{ \displaystyle y=3x+15}\) you know that \(\large\color{blue}{ \displaystyle y}\) is the same thing as \(\large\color{blue}{ \displaystyle 3x+15}\) (based on the second equation). So knowing this, you can replace \(\large\color{blue}{ \displaystyle y}\) in the first equation, with \(\large\color{blue}{ \displaystyle 3x+15}\). (this replacement is called substitution.) so lets do this "substitution". \(\large\color{blue}{ \displaystyle 2x+3y=1}\) \(\large\color{blue}{ \displaystyle 2x+3\color{red}{(3x+15)}=1}\)

OpenStudy (solomonzelman):

Now, you can solve for x.